# Resolving the Bulk-Boundary Correspondence Paradox on Low-Symmetry Surfaces of Weyl Semimetals

### v1

@@ Methods
Sample preparation of the (103) surface To obtain the low-symmetry (103) surface of NdAlSi, we first determined the crystallographic orientation of the single crystals by single crystal XRD. Based on the identified orientation, the crystals were cut with a high-precision diamond wire saw to expose the desired (103) direction. Prior to mounting, a shallow pre-cut was introduced along the target orientation, which facilitated controlled cleavage of the sample inside the ARPES chamber. This procedure enabled reliable preparation of well-defined (103) surfaces suitable for spectroscopic measurements.
@@
For the (001) surface states of NdAlSi, a slab model with 36 atomic layers and a 20 Å vacuum layer was constructed. The calculated results are consistent with experimental data. For the (103) surface states and Fermi arcs, we employed an iterative Green’s function method to model a semi-infinite surface and compute its surface states. Among the four possible terminations of the (103) surface, the Si-Al plane termination yielded results consistent with experimental observations.
@@ Acknowledgement
The work presented here was financially supported by the Knut and Alice Wallenberg foundation (2018.0104). Y.G.S. acknowledges the National Natural Science Foundation of China (Grants No. U2032204), the Informatization Plan of Chinese Academy of Sciences (CAS-WX2021SF-0102) and the National Key R&D Program of China 2024YFA1408400. Q. S. W. acknowledges the National Natural Science Foundation of China No. 12274436. We acknowledge MAX IV Laboratory for time on Beamline BLOCH under Proposal 20241395 and 20250284 as well as Diamond Light Source for time on Beamline I05 under Proposal SI39652. Research conducted at MAX IV, a Swedish national user facility, is supported by the Swedish Research council under contract 2018-07152, the Swedish Governmental Agency for Innovation Systems under contract 2018-04969, and Formas under contract 2019-02496.

### v3

@@ Methods
Sample preparation of the (103) surface To obtain the low-symmetry (103) surface of NdAlSi, we first determined the crystallographic orientation of the single crystals by single crystal XRD. Based on the identified orientation, the crystals were cut with a high-precision diamond wire saw to expose the desired (103) direction. Prior to mounting, a shallow laser pre-cut was introduced along the target orientation, which facilitated controlled cleavage of the sample inside the ARPES chamber. This procedure enabled reliable preparation of well-defined (103) surfaces suitable for spectroscopic measurements.
@@
For the (103) surface states and Fermi arcs, we employed an iterative Green’s function method to model a semi-infinite surface and compute its surface states. Among the four possible terminations of the (103) surface, the Si-Al plane termination yielded results consistent with experimental observations.
@@ Acknowledgement
The authors thank Maxim Breitkreiz for his useful discussion of bulk-boundary correspondence. The work presented here was financially supported by the Swedish Research council (2019-00701 and 2019-03486) and the Knut and Alice Wallenberg foundation (2018.0104). Y.G.S. acknowledges the National Natural Science Foundation of China (Grants No. U2032204), the Informatization Plan of Chinese Academy of Sciences (CAS-WX2021SF-0102) and the National Key R&D Program of China 2024YFA1408400. Q. S. W. acknowledges the National Natural Science Foundation of China No. 12274436. We acknowledge MAX IV Laboratory for time on Beamline BLOCH under Proposal 20241395 and 20250284 as well as Diamond Light Source for time on Beamline I05 under Proposal SI39652. Research conducted at MAX IV, a Swedish national user facility, is supported by the Swedish Research council under contract 2018-07152, the Swedish Governmental Agency for Innovation Systems under contract 2018-04969, and Formas under contract 2019-02496.